# Thermalization slowing down of weakly nonintegrable quantum spin dynamics

### v1

@@ Abstract
In the 1D spin-$1/2$ chain we study, both timescales, $T_{\lambda}$ and $T_{E}$, diverge as power-laws with decreasing integrability breaking parameter. We compare the exponents of both $T_{E}$ and $T_{\lambda}$ in the two network classes. We find that the exponents of $T_{E}$ and $T_{\lambda}$ are comparable to each other for LRN. This indicates that the rate at which thermalization (in the ETH sense) slows down upon approaching the integrable limit is comparable to the slowing down of operator spreading. We identify this as a distinguishing feature of the long-range quantum networks. The exponent for $T_{E}$ is found to be larger than $T_{\lambda}$ by at least an order of magnitude for SRN. This is indicative of the fact that ETH-like thermalization slows down exponentially faster than operator spreading as integrability is approached.
@@
We identify this drastic difference in the relative slow-down of thermalization and operator spreading as the distinguishing feature of short- and long-range networks. This is consistent with the character of thermalization slowing down in classical systems [41, 40] where classical Lyapunov time is compared to the classical ergodization time, and generalizes the notion of classical networks to quantum many-body systems.
@@
Many open questions naturally emerge from this investigation. One natural direction is testing this classification for other systems. Further, other probes might be able to distinguish and be sensitive to these two network classes. In the classical case the Lyapunov spectrum scaling close to integrability proved to be instrumental in classifying network classes [41]. It would interesting to study probes that do not have classical analogues, such as quantum entanglement [79]. Our investigation was restricted to finite system sizes, the scaling of the two timescales with system size (and therefore in the thermodynamic limit) is worth investigating. There has also been a large body of work that has investigated the exact nature of the fluctuations of operator evolution (for example, $R(t)$ in Eq. (6)). This suggests that a more concrete connection between ETH and ergodization time might exist. Random matrix theory is also expected to play a crucial role in this characterization. It would also be interesting to study the effective random matrix theory near integrability for the two network classes.
@@ I Krylov complexity
The decay of the spectral function for large $\omega$ has been the subject of intense investigation in recent years, and has been found to be a useful indicator of chaotic and integrable dynamics [78, 80, 81, 82, 55].
@@
II Passage times
@@
Figure 1: Schematic representation of time-evolution of $\langle\mathcal{O}(t)\rangle$ around the mean value $\overline{\mathcal{O}}$. We have used a log-scale on the $t$-axis.
@@ III Details of the numerics
The results for the short- and long-range networks are discussed in the main text. For the short-range network, the appropriate choice of ergodization time is the ratio of variance and mean of the excursion times as supported by the data showin in Fig. 2. The exponentially larger scale of $\sigma^{2}$ as compared to $\mu$ suggests that the fluctuations dominate the dynamics. The relative strength of the effect of the fluctuations defines an appropriate time-scale $T_{E}=\sigma^{2}(\tau)/\mu(\tau)$.
(a) Mean excursion time $\mu(\tau)$.
(b) Variance of the excursion time $\sigma^{2}(\tau)$.
Figure 2: The mean and variance of passage times for the short-range network. Here $\tau_{\pm}$ denotes positive and negative excursion times.
For the long-range network, the mean and variance are of equivalent orders. This suggests that the dynamics is proportionately dominated by the averages as well as the fluctuations. Therefore, one can pick either the mean or variance as the relevant ergodization time $T_{E}$. We choose the mean as the ergodization time $T_{E}=\mu(\tau)$. This choice is supported by the data shown in Fig. 3. We see that the ratios of the mean and the variance remain of the same order as the integrability breaking parameter is varied. Therefore they do not represent useful timescales that are sensitive to the integrability breaking effect.
(a) Ratio of variance and mean $\sigma^{2}(\tau)/\mu(\tau)$.
(b) Ratio of standard deviation and mean $\sigma(\tau)/\mu(\tau)$.
Figure 3: The ration (a) variance and mean and (b) standard deviation and mean of passage times for the long-range network. Here $\tau_{\pm}$ denote positive and negative excursion times.

### v2

@@ Abstract
In the 1D spin-$1/2$ chain we study, both timescales, $T_{\lambda}$ and $T_{E}$, diverge as power-laws with decreasing integrability breaking parameter. We compare the exponents of both $T_{E}$ and $T_{\lambda}$ in the two network classes. We find that the exponents of $T_{E}$ and $T_{\lambda}$ are comparable to each other for LRN and SRN. However, this result uses varying different model parameters. In order to be able to quantitatively compare the relative growth of the time scales for both regimes, we measure $T_{E}$ in units of $T_{\lambda}$ and plot the outcome as a function of $T_{\lambda}$. We then find that the LRN regime shows a relatively small rate at which thermalization (in the ETH sense) slows down upon approaching the integrable limit compared to operator spreading. For the SRN case instead, the rate at which thermalization (in the ETH sense) slows down upon approaching the integrable limit compared to operator spreading, is much larger. Therefore, in the SRN case it takes more and more time to thermalize as compared to the operator growth time scales.
This implies that in the SRN regime, ETH-like thermalization slows down exponentially faster than operator spreading as integrability is approached.
@@
We identify this drastic difference in the relative slow-down of thermalization and operator spreading as the universal feature of short-range networks, as opposed to long-range networks. This is similar to the character of thermalization slowing down in classical systems [41, 40] where a classical Lyapunov time is compared to the classical ergodization time, and the two network classes respond in very different manners.
@@
Many open questions naturally emerge from this investigation. One natural direction is testing this classification for other types of quantum systems. Further, other probes might be able to distinguish and be sensitive to these two network classes. In the classical case, the Lyapunov spectrum scaling close to integrability proved to be instrumental in classifying network classes [41]. It would interesting to study probes that do not have classical analogues, such as quantum entanglement [83], in such phenomena. Our investigation was restricted to finite system sizes, the scaling of the two timescales with system size (and therefore in the thermodynamic limit) is also worth investigating. There has also been a large body of work that has investigated the exact nature of the fluctuations of operator evolution (for example, $R(t)$ in Eq. (6)). This suggests that a more concrete connection between ETH and ergodization time might exist. Random matrix theory is also expected to play a crucial role in this characterization. It would also be interesting to study the effective random matrix theory near integrability for the two network classes.
@@ I Krylov complexity
The decay of the spectral function for large $\omega$ has been the subject of intense investigation in recent years and has been found to be a useful indicator of chaotic and integrable dynamics [81, 85, 86, 87, 55].
@@
II Conserved Quantities: SRN
When studying the SRN limit and computing the thermalization timescales, we choose the conserved quantity (in the integrable limit $h\rightarrow 0$) to be the Pauli matrix $\sigma^{z}_{i}$, at some lattice site $i$. This quantity is local, and so the integrability breaking induces extra terms that are also local (but not necessarily $1-$ local). It is interesting to consider what would happen if a non-local conserved quantity is instead considered in the SRN case. The results do not change much for the following reasons: Let us consider the non-local initial operator
$\mathcal{O}{(N)}=\sum_{i=1}^{N}\sigma^{z}_{i}$ (6)
where the superscript $(N)$ is used to indicate that the operator is non-local. Correspondingly, the local operator is denoted as $\mathcal{O}^{(1)}=\sigma^{z}_{1}$. The time evolution of the operator can be broken into that of individual $\sigma^{z}_{i}$. The evolution for each of these should be equivalent since the state (with respect to which the expectation value is calculated) |$\ket{\psi}$ is a random state and hence has roughly equal weight at each site $i$. This allows us to approximate the fluctuation equation as
$\langle\mathcal{O}^{(N)}(t)\rangle-\overline{\mathcal{O}^{(N)}}\approx N\left(\langle\mathcal{O}^{(1)}(t)\rangle-\overline{\mathcal{O}^{(1)}}\right)\,,$ (7)
which has a similar distribution of zeros, and hence the similar moments, as that of $\mathcal{O}^{(1)}$.
The Krylov complexity (or Lanczos growth) of such non-local operators should also be the same as that of local operators. This is because we are working with translation symmetric systems, and therefore each $\sigma_{i}$ can be replaced by $\sigma_{1}$ in the BCH expansion. The resulting overall factor of $N$ is taken care of via normalization.
To support this argument, we present the numerical results for the two cases in Fig. 1. The non-local timescales are represented by a superscript $(N)$.
Figure 1: Comparison of ergodization times and Lyapunov times for SRN with local and non-local operators. The superscript $(N)$ stands for the non-local operator. The initial states are different but chosen from the same random distribution, for the local and non-local cases.
It is interesting to note that the ergodization timescales for $\tau_{\pm}$ are much more similar for the case of $\mathcal{O}^{(N)}$ than for $\mathcal{O}^{(1)}$. This is due to the sum of local operators having a “smoothing” effect on the random state |$\ket{\psi}$. This causes the timescales obtained from $\tau_{+}$ and $\tau_{-}$ to almost exactly overlap in the weakly integrable limit.
III Conserved Quantities: LRN
Here we discuss the conserved quantities of the integrable limit $g\rightarrow 0$ of the long-range network class. This integrable Hamiltonian is known as the transverse field Ising model and has a complete set of conserved quantities. These are given by
$I^{(k)}$ $=iJ\sum_{j=1}^{N}\left(S^{zy}_{j:j+k}-S^{yz}_{j:j+k}\right)\;\;\;\;k=1,\dots,N-1$ (8)
where we have the following shorthand
$S^{\alpha\beta}_{j:j+l}=\sigma^{\alpha}_{j}\left(\prod_{n=1}^{l-1}\sigma^{x}_{j+n}\right)\sigma^{\beta}_{j+l}$ (9)
It is straightforward to see that $[H_{\text{TFIM}},I^{(k)}]=0$. These quantities can be interpreted as linear combinations of mode occupation numbers in the Jordan-Wigner fermion theory [89]. For $k=N$, the conserved quantity corresponds to $\prod_{i=1}^{N}\sigma^{x}_{i}$, which is a symmetry operation corresponding to the replacement $\sigma^{z,y}\rightarrow-\sigma^{z,y}$.
Figure 2: Scaling of ergodization times for LRN (for all conserved quantities $I^{(k)}$) and comparison with the same for SRN (single operator).
We compare the ergodization times $T_{E,\pm}$ for each $I_{k}$, with the respective Lyapunov time $T_{\lambda}$ obtained via the Krylov method. The results are presented in Fig. 2. It is evident that the scaling is different for different $I^{(k)}$. Note that despite different scaling behaviour of the ergodization time $T_{E}$ for different conserved quantities $I^{(k)}$, the ergodization time for the $\sigma^{z}$ operator in the SRN case bounds the $T_{E}$ for all $I^{(k)}$ from above. Assuming a power law scaling behavior for the timescales
$\log T_{E,\lambda}=\begin{cases}&\alpha_{E,\lambda}\log g+\delta_{E,\lambda}\;\;,\;\;\text{LRN}\\ &\alpha_{E,\lambda}\log h+\delta_{E,\lambda}\;\;,\;\;\text{SRN}\end{cases}$ (10)
we obtain the exponents listed in Table 1.
$k$ $\alpha^{+}_{E}$ $\delta^{+}_{E}$ $\alpha_{\lambda}$ $\delta_{\lambda}$
$1$ $-1.65828$ $-0.395546$ $-0.046659$ $0.262234$
$2$ $-1.28645$ $-0.592328$ $-0.0204276$ $0.248453$
$3$ $-1.9492$ $-0.638931$ $-0.151105$ $0.00884637$
$4$ $-0.734752$ $-0.356261$ $-0.169417$ $-0.0685405$
$5$ $-3.03973$ $-2.60729$ $-0.154005$ $-0.0406314$
$6$ $-1.54398$ $-0.570509$ $-0.155603$ $-0.0419308$
$7$ $-2.49685$ $-1.76273$ $-0.196703$ $-0.11035$
Table 1: Power-law coefficients for all $I^{(k)}$ for the LRN case. The superscript $+$ stands for results extracted from positive passage times. The results for negative passage times are comparable.
For the corresponding SRN case, we observe that the coefficients are
$\alpha^{+}_{E,SRN}=-2.26301\;\;,\;\;\delta^{+}_{E,SRN}=-0.136552$
$\alpha_{\lambda,SRN}=-0.0817788\;\;,\;\;\delta_{\lambda,SRN}=0.217633$
It is also instructive to compare the scaling of the ratio of timescales $T_{E}/T_{\lambda}$ as a function of $T_{\lambda}$. This ratio is expected to diverge upon approaching the integrable limit (i.e. as $T_{\lambda}$ increases). Since different conserved quantities will in general have different ranges of values of $T_{\lambda}$ (although within the same order), it is better to instead study the ratio as a function of $T_{\lambda}/\text{max}(T_{\lambda})$. We present this result in Fig. 3.
Figure 3: Behavior of the ratio $T_{E}/T_{\lambda}$ as a function of $T_{\lambda}/T^{*}_{\lambda}$, where $T^{*}_{\lambda}=\text{max}(T_{\lambda})$. The scaling for all $I^{(k)}$ (LRN) is compared to that of $\sigma^{z}$ (SRN). These results are presented for the integrability-breaking parameter close to 00.
Finally, we consider the scaling of $T_{\lambda}$ and $T_{E}/T_{\lambda}$ with the integrability-breaking parameter $g,h$ for completeness. This is presented in Fig. 4 and Fig. 5 respectively.
Figure 4: Behavior of the Lyapunov times $T_{\lambda}$ as a function of the integrability breaking parameter $g,h$. The region close to the integrable limit is explored. It is observed that for all conserved quantities $I^{(k)}$ (LRN), as well as for the SRN observable $\sigma^{z}$, the Lyapunov times are comparable.
Figure 5: Behavior of the ratio $T_{E}/T_{\lambda}$ as a function of the integrability breaking parameter $g,h$. The results support the conclusion of Fig. 2.
The results of Fig. 4 and 2 explain the observation in Fig. 5, since the Lyapnuov times $T_{\lambda}$ for all $I^{(k)}$ (LRN) and $\sigma^{z}$ (SRN) remain comparable throughout the range of $g,h$ explored. However, the ergodization times scale in a different manner (significantly different, on the log-scale, as seen in Fig.  2). Thus the ratio $T_{E}/T_{\lambda}$ is also highly sensitive to initial operator choice and the universality classification.
IV Passage times
@@
Figure 6: Schematic representation of time-evolution of $\langle\mathcal{O}(t)\rangle$ around the mean value $\overline{\mathcal{O}}$. We have used a log scale on the $t$-axis.
@@ V Details of the numerics
The results for the short- and long-range networks are discussed in the main text. For the short- and long-range network, the appropriate choice of ergodization time is the ratio of variance and mean of the excursion times as supported by the data shown in Fig. 7 and Fig. 8. The exponentially larger scale of $\sigma^{2}$ as compared to $\mu$ suggests that the fluctuations dominate the dynamics. Therefore, the appropriate choice of a timescale would be a ratio of the fluctuation to the mean, given by $\frac{\sigma^{2}}{\mu}$. This is found to be the case in both LRN and SRN.
(a) Mean of ergodization times $\mu(\tau)$.
(b) Variance of ergodization times $\sigma^{2}(\tau)$.
Figure 7: Mean and Variance of ergodization times for all operators $I^{(k)}$ and the SRN operator $\sigma^{z}$, as a function of $g$ and $h$ respectively.
(a) Mean of ergodization times $\mu(\tau)$.
(b) Variance of ergodization times $\sigma^{2}(\tau)$.
Figure 8: Mean and Variance of ergodization times for the averaged operator $N^{-1}\sum_{k}I^{(k)}$ and the SRN operator $\sigma^{z}$, as a function of $g$ and $h$ respectively.